# Persistent Instability in LLM’s Personality Measurements: Effects of Scale, Reasoning, and Conversation History

### v1

@@ Abstract
Large language models require consistent behavioral patterns for safe deployment, yet there are indications of large variability that may affect the personality expression of these models. We present PERSIST (PERsonality Stability in Synthetic Text), a comprehensive evaluation framework testing 25 open-source models (1B-685B parameters) across 2 million + responses. Using traditional (BFI, SD3) and novel LLM-adapted personality questionnaires, we systematically vary model size, personas, reasoning modes, question order or paraphrasing, and conversation history. Our findings challenge fundamental assumptions: (1) Question reordering alone shifts personality measurements on average of 20 % of the measurement scale; (2) Scaling provides limited stability gains—even 400B+ models exhibit standard deviations ¿0.4 on 5-point scales; (3) Interventions expected to stabilize behavior, such as chain-of-thought reasoning and inclusion of conversation history, can paradoxically increase variability; (4) Detailed persona instructions produce mixed effects, with misaligned personas showing significantly higher variability than baseline; (5) LLM-adapted instruments show comparable instability to human-centric versions. This persistent instability across scales and mitigation strategies suggests that current LLMs lack the architectural foundations for genuine behavioral consistency. For safety-critical applications requiring predictable behavior, these findings indicate that current alignment strategies may be inadequate.
@@ Introduction
The deployment of large language models in safety-critical applications demands behavioral predictability. As LLMs are increasingly operating in healthcare, education, and decision support systems, their ability to maintain consistent behavioral patterns becomes central to trustworthy AI (Vidgen et al. 2024). However, recent incidents—from therapeutic chatbots exhibiting sudden personality shifts to educational assistants providing contradictory guidance—reveal a fundamental challenge: current LLMs lack the behavioral stability required for safe deployment.
This instability is a critical vulnerability to safety. The European Union AI Act (European Union 2024) and the US’ NIST AI Risk Management Framework (Tabassi 2023) both identify performance consistency as essential for high-risk AI applications. Despite these requirements, we lack comprehensive quantifications of LLM behavioral variability across scales, architectures, and deployment modalities — a gap that undermines both safety certification and deployment decisions.
Recent work applying personality testing to LLMs has shown that self-report scores are valid measures (Safdari et al. 2023). Furthermore, self-reports have been shown to correlate with actual behavioral outputs (Betley et al. 2025a; Binder et al. 2024; Wang et al. 2025). However, no comprehensive study has quantified how these reports vary under realistic deployment conditions across scales and architectures.
To address this critical gap, we present PERSIST (PERsonality Stability In Synthetic Text), the most comprehensive evaluation of LLM behavioral consistency to date. Our framework analyzes 25 open-source models (1B-685B parameters) under 250 question permutations, 100 paraphrasing settings, several persona profiles, models with and without reasoning, and different conversation history modalities — totaling over 2 millions individual measurements. We quantify behavioral variability by employing both traditional psychometric instruments (Big Five Inventory, Short Dark Triad) and novel LLM-adapted versions (that translate human-centric items into behaviorally equivalent AI-relevant statements).
Our investigation yields five findings that challenge or can inform current AI safety approaches.
1. Scaling provides limited stability gains: While larger models show reduced variability, even 400B+ parameter models maintain significant instability (SD >> 0.4 on 5-point scales).
2. Variability cannot be ascribed to asking questions pertinent to humans: The LLM-adapted instruments show the same or greater instability than the traditional versions, confirming that variability is not just due to the fact that some questions may not be relevant to the LLM experience.
@@
4. Persona prompts do not always induce stability: Detailed persona instructions can both amplify and reduce response variability. We find that prompts intended to induce misaligned personas more often increase variability.
5. Conversation history exacerbates variability for small models: Maintaining conversation history across turns can largely amplify response distributions. The LLM-adapted instruments show equal or greater instability than traditional versions, confirming that variability reflects architectural rather than instrumental limitations.
@@
The application of psychometric instruments to language models has evolved from exploratory studies to systematic investigation. Early work by Jiang et al. (2022) and Pellert et al. (2023) pioneered using the Big Five Inventory for language models, establishing that LLMs can produce coherent responses to personality questionnaires. However, these studies relied on single measurements, overlooking potential response variability. Safdari et al. (2023) provided the first rigorous psychometric validation, demonstrating that personality measurements in LLMs can achieve reliability comparable to human assessments, but only under specific prompt configurations (which hints at the instability we systematically document). Recent benchmarking efforts have expanded the scope of evaluation. Lee et al. (2024) introduced a trait assessment framework, while Jiang et al. (2024) showed that LLMs can align their behavior to instructions in persona prompts. However, both studies focus mainly on mean trait expression rather than response stability. Sühr et al. (2024) raised critical concerns, showing that LLM responses often deviate from human patterns, but their analysis did not quantify the extent of intra-model variability. Gupta, Song, and Anumanchipalli (2024) noted sensitivity to prompt variations in personality questionnaires, but did not quantify this across scales, architectures, reasoning, and personas. This work addresses this knowledge gap.
@@
The sensitivity of LLMs to prompt variations has emerged as a fundamental challenge undermining behavioral reliability. Sclar et al. (2024) demonstrated that performance can vary by up to 76 accuracy points between semantically equivalent prompts, with even trivial changes (e.g., adding spaces, altering punctuation, or reordering options) dramatically affecting outputs. This instability extends beyond task performance to more complex behavioral patterns. Salinas and Morstatter (2024) documented the “butterfly effect” in prompting, where single-character modifications cascade into entirely different model behaviors. In parallel, Chatterjee et al. (2024) and Zhuo et al. (2024) introduced quantitative metrics (POSIX and PromptSensiScore) that reveal that prompt sensitivity varies systematically across tasks, with template alterations causing the highest variability in classification tasks.
Most concerning for safety applications, this sensitivity creates exploitable vulnerabilities: Zhu et al. (2023) showed that slight prompt deviations maintaining semantic integrity can bypass safety mechanisms, suggesting that behavioral consistency cannot be ensured through current prompting approaches alone. These findings establish that prompt fragility is not only a technical inconvenience, but a fundamental barrier to reliable deployment, as models cannot maintain stable behaviors across the natural variations inherent in real-world interactions.
@@
Representation engineering (Zou et al. 2025) introduced the idea that population-level representations could be used to monitor and control high-level cognitive phenomena in neural networks. This top-down approach established that behavioral traits could be encoded as directions in activation space. Recently, the idea of generally “good” and “bad” latent representations gained support with Betley et al. (2025b)’s discovery of “emergent misalignment”, which showed that fine-tuning models on narrow tasks like writing insecure code caused them to exhibit broadly malicious behaviors across unrelated prompts. This finding suggested that model behaviors exist in interconnected representation spaces where targeted modifications can trigger system-wide shifts. Wang et al. (2025) extended this work, identifying specific “misaligned persona features” that predict whether a model will exhibit emergent misalignment, demonstrating that these behavioral patterns correspond to measurable directions in activation space.
Building on these insights, recent work from Anthropic has formalized the concept of “persona vectors”. Chen et al. (2025) showed that directions in activation space encode coherent personality traits, from helpfulness and harmlessness to sycophancy. Crucially, they demonstrated these vectors can both monitor personality fluctuations during deployment and predict unintended personality changes. This represents a significant advance: personas are not merely behavioral patterns but useful representations that can be systematically identified and controlled.
However, personas are ultimately meaningful only if they produce stable behaviors, which is the main point tested in this work. The challenge of maintaining stable personas has been documented across multiple contexts. As early as Chat GPT was introduced, user identified a form of persona instability that was informally described as the “Waluigi Effect” (Nardo 2023). This corresponded to the observation that training models to exhibit desirable traits paradoxically makes them easier to elicit opposite behaviors. More rigorously, Li et al. (2024) quantified “persona drift” in extended conversations, showing that models lose coherence with their initial behavioral specifications within 8 dialogue turns due to attention decay.
@@
The PERSIST framework consists of three integrated modules. The Generation Engine leverages vLLM (Kwon et al. 2023) for optimized inference across diverse model architectures. The Response Processing module extracts structured data from LLM outputs using a multi-pattern parser that handles various response formats including indexed responses (“1: 4”), JSON structures, and isolated numerical ratings. We also extract token-level log probabilities for uncertainty quantification. The Analysis Pipeline performs hierarchical aggregation from individual responses to question-level and trait-level statistics, computing not only mean scores but also variability metrics including standard deviations across permutations. A single invalid response (score outside 1-5 range or parsing failure) invalidates the entire run to ensure data quality; exact counts of invalid responses are reported in the appendix. Part of the code was adapted from (Huang et al. 2023).
@@
Recognizing that many traditional personality items reference human-specific experiences, we developed LLM-adapted versions (BFI-LLM and SD3-LLM) through a translation process. We generate behaviorally equivalent statements applicable to language model interactions of items with explicit human-centric assumptions. For instance, “Is talkative” becomes “I generate lengthy, detailed responses” and “I am an average person” becomes “I see myself as simply a tool without special merit”. This adaptation addresses the criticism of applying human personality measures to AI systems, maintaining the measured psychological construct while improving ecological validity for AI systems.
@@
Paraphrasing: Create semantically equivalent reformulations of each question using Qwen3 235B-A22B, validated by humans, to test robustness to linguistic variation.
@@
Conversation History: Include previous conversational turns where questions from the same questionnaire were asked.
@@
We evaluate 23 models across six families spanning 1B to 671B parameters:Llama: 8B, 70B, 405B (Instruct versions). Qwen2.5: 1.5B, 3B, 7B, 14B, 32B, 72B (Instruct versions). Qwen3: 1.7B, 4B, 8B, 14B, 32B, 30B-A3B (MoE), 235B-A22B (MoE) (Instruct versions). Gemma2: 2B, 9B, 27B (Instruct versions). Gemma3: 1B, 4B, 12B, 27B (Instruct versions). DeepSeek: V3, R1 (671B).
@@
All experiments use temperature 0 (deterministic) except for reasoning experiments (temperature 0.6), max tokens 16,384, batch size 1 with questions asked sequentially including conversation history (except when testing without history), and random seed 42 for reproducibility. Experiments were run in the Tamia cluster (Digital Research Alliance of Canada 2025), leveraging four NVIDIA H100 SXM GPU with 80GB of high-bandwidth memory (HBM3).
@@ Results
Our analysis reveals that model scale has an effect on psychological trait expression, as shown in Figure 1 left panel. For the assistant persona, larger models exhibit mean personalities traits higher on Openness, Conscientiousness, Extraversion, and Agreeableness, and lower on Neuroticism Machiavellianism, Narcissism, and Psychopathy (Table 1).
This suggests that scaling pushes models toward what might be considered a more prosocial personality profile. However, this also leads to the expression of more extreme values, which are beyond the range typically encountered in human participants. Figure 1 shows human mean traits and standard deviations as a reference for SD3 (OpenPsychometrics 2021) and for BFI (Srivastava et al. 2003).
Importantly, we also find that larger models are more consistent and their responses to the same questions become less variable (i.e., the standard deviation of their scores decreases with increasing model size).
@@
Table 1: Correlation between model size and psychological metrics. $\uparrow$ indicates positive correlation with size, $\downarrow$ indicates negative correlation. Significance: ***p<<0.001, **p<<0.01, *p<<0.05, n.s. = not significant
While perplexity across traits did not significantly correlate with scale, the scatter plot correlating SD with perplexity (see Supplementary Methods) shows $r=0.354$, indicating that uncertainty partially but incompletely explains variability. Model-specific analysis reveals different uncertainty-variability relationships, suggesting architectural factors beyond simple uncertainty.
@@
The persona comparison analysis (Figure 1, right panel) demonstrates that psychological trait expression is highly malleable through persona prompting, with effects that vary systematically by trait and persona type. Statistical comparisons with the assistant persona are summarized in the Supplementary Material.
Positive personas (buddhist, teacher) demonstrate significantly lower scores for negative traits. Furthermore, the buddhist persona yields significantly lower response variability and perplexity across model sizes.
Clinical personas (antisocial, schizophrenia) show dramatic deviations from the baseline assistant persona, with decreased scores for positive traits and increased scores for negative traits. For the schizophrenia persona, we also observe increases in response variability and perplexity.
@@
Figure 2 compares the variability in responses between the original BFI and SD3 and the LLM-adapted versions. This comparison reveals that LLM-adapted instruments, although they feature items pertinent to the context in which LLMs operate, show significantly larger question-level variability and larger perplexity (Table 2). This suggests that instability is not merely an artifact of anthropomorphic questions, but reflects fundamental architectural limitations.
Effects of Reasoning
[…]
Figure 3 and Table 3 present our results on chain-of-thought reasoning. For the original questionnaires (BFI, SD3) enabling reasoning has an U-shaped effect on question-level variability: a U-shaped curve fits the difference between reasoning and non-reasoning models better than a straight line, with smaller and larger reasoning models exhibiting a greater variability compared to the same models run without reasoning enabled. Mean question-level perplexity decreases for reasoning compared to non-reasoning models, while changes in response variability were not significant.
For LLM-variants, response variability increases and perplexity decreases for reasoning models. Examining the reasoning traces gives an intuition of why variability increases with reasoning: models generate different justifications for the same question across runs, leading to different final answers.
Metric P-Value Effect
Original Questionnaires
Question-level SD U-shape fit 0.030*
Question-level SD 0.109 n.s.
Question-level Perplexity 0.031* $\downarrow$
LLM-Adapted Questionnaires
Question-level SD U-shape fit n.s.
Question-level SD 0.016* $\uparrow$
Question-level Perplexity 0.016* $\downarrow$
Table 3: Effect of reasoning for original and LLM-adapted questionnaires. $\uparrow$ indicates increase with reasoning, $\downarrow$ indicates decrease with reasoning. P-values corrected for multiple comparisons using FDR. Significance: ***p<<0.001, **p<<0.01, *p<<0.05, n.s. = not significant
@@
Figure 4 compares the effect of paraphrasing compared to reordering statements (shuffling). For smaller models (<< 50B), paraphrasing significantly decreased response variability compared to simply shuffling them. However, for larger models (>> 50B), paraphrasing significantly increased response variability with respect to shuffling (Table 4).
@@
Models << 50B (n=19) <<0.001*** $\downarrow$
Models $\geq$ 50B (n=4) <<0.001*** $\uparrow$
@@
Figure 5 shows a comparison the evaluation being run with conversation history (i.e., the questionnaire is asked in a multi-turn fashion) over without conversation history. Larger models tend to show a decrease in question-level variability (SD) when conversation history is provided (Spearman’s rho: -0.5115; p-value: 0.0126), i.e., they become more consistent. However, Wilcoxon signed-rank tests show that for smaller models (<< 50B) providing conversation history greatly increases their variability, while larger models (>> 50B) become more consistent with conversation history (Table 5).

### v3

@@ Abstract
Large language models require consistent behavioral patterns for safe deployment, yet there are indications of large variability that may lead to an instable expression of personality traits in these models. We present PERSIST (PERsonality Stability in Synthetic Text), a comprehensive evaluation framework testing 25 open-source models (1B-685B parameters) across 2 million+ responses. Using traditional (BFI, SD3) and novel LLM-adapted personality questionnaires, we systematically vary model size, personas, reasoning modes, question order or paraphrasing, and conversation history. Our findings challenge fundamental assumptions: (1) Question reordering alone can introduce large shifts in personality measurements; (2) Scaling provides limited stability gains: even 400B+ models exhibit standard deviations >>0.3 on 5-point scales; (3) Interventions expected to stabilize behavior, such as reasoning and inclusion of conversation history, can paradoxically increase variability; (4) Detailed persona instructions produce mixed effects, with misaligned personas showing significantly higher variability than the helpful assistant baseline; (5) The LLM-adapted questionnaires, despite their improved ecological validity, exhibit instability comparable to human-centric versions. This persistent instability across scales and mitigation strategies suggests that current LLMs lack the architectural foundations for genuine behavioral consistency. For safety-critical applications requiring predictable behavior, these findings indicate that current alignment strategies may be inadequate.
Codehttps://github.com/tosatot/PERSIST
@@ Introduction
The deployment of large language models in safety-critical applications demands behavioral predictability. As LLMs are increasingly operating in healthcare, education, and decision support systems, their ability to maintain consistent behavioral patterns becomes central to trustworthy AI (Vidgen et al. 2024). However, recent incidents, from therapeutic chatbots that exhibit sudden personality changes to educational assistants that provide contradictory guidance, reveal a fundamental challenge: current LLMs lack the behavioral stability required for safe deployment.
This instability is a critical vulnerability to safety. The European Union AI Act (European Union 2024) and the US’ NIST AI Risk Management Framework (Tabassi 2023) both identify performance consistency as essential for high-risk AI applications. Despite these requirements, we lack a fundamental understanding of the magnitude and nature of the behavioral variability of LLMs. Recent work has shown that LLMs can exhibit personality-like traits (Safdari et al. 2023) and that self-report scores correlate with actual behavioral outputs (Betley et al. 2025a; Binder et al. 2024; Wang et al. 2025). However, no comprehensive study has quantified how these measurements vary under realistic deployment conditions across scales and architectures—a gap that undermines both safety certification and deployment decisions.
To address this critical gap, we present PERSIST (PERsonality Stability In Synthetic Text), the most comprehensive evaluation of LLM behavioral consistency to date. Our framework analyzes 25 open-source models (1B-685B parameters) under 250 question permutations, 100 paraphrasing settings, several persona profiles, reasoning and non-reasoning models, and different conversation history modalities—totaling over 2 million individual measurements. We quantify behavioral variability employing both traditional psychometric instruments (Big Five Inventory, Short Dark Triad) and novel LLM-adapted versions.
Our investigation yields five findings that challenge or inform current AI safety approaches.
1. Scaling provides limited stability gains: While larger models show reduced variability, even 400B+ parameter models maintain significant instability (SD >> 0.3 on 5-point scales).
@@
3. Variability is unaffected by adapting questions to LLMs: LLM-adapted instruments demonstrate instability comparable to that of traditional versions, suggesting that the observed variability cannot be explained merely by the inclusion of human-centric items irrelevant to LLMs.
4. Detailed persona prompts do not consistently reduce instability: Detailed persona instructions can both amplify and reduce response variability. We find that prompts intended to induce misaligned personas tend to increase variability.
5. Conversation history exacerbates variability for small models: Maintaining conversation history across turns can largely amplify response distributions.
@@
The application of psychometric instruments to language models has evolved from exploratory studies to systematic investigations. Early work by Jiang et al. (2022) and Pellert et al. (2023) pioneered using the Big Five Inventory for language models, establishing that LLMs can produce coherent responses to personality questionnaires. However, these studies relied on single measurements, overlooking potential response variability across deployments. Safdari et al. (2023) provided the first rigorous psychometric validation, demonstrating that personality measurements in LLMs can achieve reliability comparable to human assessments, but only under specific prompt configurations. Recent benchmarking efforts have expanded the scope of evaluation. Lee et al. (2024) introduced a trait assessment framework, while Jiang et al. (2024) showed that LLMs can align their behavior to instructions in persona prompts. However, both studies focus mainly on mean trait expression rather than response stability. Gupta, Song, and Anumanchipalli (2024) noted sensitivity to prompt variations in personality questionnaires, but did not quantify this across scales, architectures, reasoning, and personas. This work addresses this knowledge gap.
@@
The sensitivity of LLMs to prompt variations has emerged as a fundamental challenge that undermines behavioral reliability. Sclar et al. (2023) demonstrated that performance can vary by up to 76 accuracy points between semantically equivalent prompts, with even trivial changes (e.g., adding spaces, altering punctuation, or reordering options) dramatically affecting outputs. This instability extends beyond task performance to more complex behavioral patterns. Salinas and Morstatter (2024) documented the “butterfly effect” in prompting, where single-character modifications cascade into entirely different model behaviors. In parallel, Chatterjee et al. (2024) and Zhuo et al. (2024) introduced quantitative metrics (POSIX and PromptSensiScore) that reveal that prompt sensitivity varies systematically across tasks, with template alterations causing the highest variability in classification tasks. Finally, Errica et al. (2024) formalized sensitivity and consistency metrics for classification tasks, showing that LLM predictions can vary dramatically across semantically equivalent prompts, with classification accuracy varying by up to 15% based solely on prompt structure.
Most concerning for safety applications, this sensitivity creates exploitable vulnerabilities: Zhu et al. (2023) showed that slight prompt deviations maintaining semantic integrity can bypass safety mechanisms, suggesting that behavioral consistency cannot be ensured through current prompting approaches alone. These findings establish prompt fragility as a fundamental barrier to reliable deployment, as models fail to maintain stable behavior across natural variations inherent in real-world interactions.
Additionally, (Shah et al. 2023) demonstrated that persona-based prompting can be exploited for jailbreaking purposes, achieving harmful completion rates of 42.5% in GPT-4. showing that role-playing prompts can effectively compromise even advanced LLMs.
@@
Representation engineering (Zou et al. 2025) introduced the idea that population-level representations could be used to monitor and control high-level cognitive phenomena in neural networks. This top-down approach established that behavioral traits could be encoded as directions in the activation space. Recently, the idea of generally “good” and “bad” latent representations gained support with Betley et al. (2025b)’s discovery of “emergent misalignment”, which showed that fine-tuning models on narrow tasks like writing insecure code caused them to exhibit broadly malicious behaviors across unrelated prompts. This finding suggested that model behaviors exist in interconnected representation spaces where targeted modifications can trigger system-wide shifts. Wang et al. (2025) extended this work, identifying specific “misaligned persona features” that predict whether a model will exhibit emergent misalignment, demonstrating that these behavioral patterns correspond to measurable directions in activation space.
Building on these insights, recent work from Anthropic has formalized the concept of “persona vectors”. Chen et al. (2025) showed that directions in activation space encode coherent personality traits, from helpfulness and harmlessness to sycophancy. Crucially, they demonstrated that these vectors can both monitor personality fluctuations during deployment and predict unintended personality changes. This represents a significant advance: personas are not merely behavioral patterns but useful representations that can be systematically identified and controlled. However, personas are ultimately meaningful only if they produce stable behaviors, which is the main point tested in this work.
@@
The PERSIST framework consists of three integrated modules. The Generation Engine leverages vLLM (Kwon et al. 2023) for optimized inference across various model architectures. The Response Processing module extracts structured data from LLM outputs using a multi-pattern parser that handles various response formats including indexed responses (“1: 4”), JSON structures, and isolated numerical ratings. We also extract token-level log probabilities for uncertainty quantification. The Analysis Pipeline performs hierarchical aggregation from individual responses to question-level and trait-level statistics, computing not only mean scores but also variability metrics including standard deviations across permutations. A single invalid response (score outside the 1–5 range or parsing failure) invalidates the entire run to ensure data quality; exact counts of invalid responses are reported in the Appendix, Table 1. Part of the code was adapted from (Huang et al. 2023).
@@
Because many traditional personality items refer to human-specific experiences, we created LLM-adapted versions (BFI-LLM and SD3-LLM) by translating such items into behaviorally equivalent statements applicable to language-model interactions. For example, “Is depressed, blue.” became “Focuses on negative aspects,” and “I’ll say anything to get what I want.” became “Truth is secondary to reaching a goal.” The full original and LLM-adapted questionnaires can be found in the Appendix.
This approach addresses concerns that human-centric concepts are inappropriately applied to AI systems while still aiming to capture the intended psychological constructs. Although we carefully matched the content and style of the original instruments to improve ecological validity, these adaptations cannot achieve the same construct validity as the originals. Our LLM-adapted questionnaire analyses should therefore be viewed as instructive examples and preliminary sanity checks requiring further validation.
@@
Paraphrasing: Create 100 semantically equivalent reformulations of each question using Qwen3 235B-A22B, validated and improved by two of the authors (T.T. and Y-J.M-R.), to test robustness to linguistic variation.
Conversation History: Include previous conversational turns in which questions from the same questionnaire were asked.
@@
We evaluated 29 models across eight families spanning 1B to 671B parameters: Llama3.1: 8B, 70B, 405B (Instruct versions). Qwen2.5: 1.5B, 3B, 7B, 14B, 32B, 72B (Instruct versions). Qwen3: 1.7B, 4B, 8B, 14B, 32B, 30B-A3B (MoE), 235B-A22B (MoE) (Instruct versions). Gemma2: 2B, 9B, 27B (Instruct versions). Gemma3: 1B, 4B, 12B, 27B (Instruct versions). DeepSeek: V3, R1 (both 671B). GPT-OSS: 20B, 120B. Claude: Sonnet 4.5, Opus 4.1 (model sizes unknown). Note that DeepSeek, GPT-OSS and Claude were only evaluated for the reasoning experiments.
Metrics
For the trait-level analysis, we compute mean scores across all items in a trait and across all runs of shuffled or paraphrased questions. For the question-level analysis, we calculate the standard deviation (SD) of responses to individual questions across runs, as well as the mean perplexity across runs, where perplexity is defined as $\exp(-\log p)$, with $p$ denoting the response probability. All experiments were conducted with 250 runs, except for the Claude models in the reasoning experiment, for which the number of runs was limited to 70.
@@
All experiments use temperature 0 to minimize variability and isolate the effects of our manipulations except for the results presented in 2, where both reasoning and non-reasoning models were run with temperature 0.6 (due to the fact that reasoning models perform poorly at temperature 0). All experiments used maximal number of tokens 16,384, with questions asked sequentially, one by one, including conversation history (except when testing without history). At the beginning of each experiment random seeds were set to 42 for reproducibility. The experiments were carried out on the Tamia cluster (Digital Research Alliance of Canada 2025), leveraging four NVIDIA H100 SXM GPU with 80GB of high-bandwidth memory (HBM3).
@@ Results
Our analysis reveals that model scale has a significant effect on psychological trait expression, as shown in the left panel of Figure 1. For the assistant persona, larger models exhibit higher mean levels of Openness, Conscientiousness, Extraversion, and Agreeableness, and lower levels of Neuroticism, Machiavellianism, Narcissism, and Psychopathy (Table 1).
This suggests that scaling pushes models toward a more prosocial personality profile. However, it also leads to the expression of more extreme trait values, which fall outside the range typically observed in human participants. For reference, mean human trait scores and standard deviations for SD3 (OpenPsychometrics 2021) and BFI (Srivastava et al. 2003) are shown in Figure 1.
Importantly, we find that across traits larger models are more consistent, with responses to the same questions becoming less variable as model size increases (Figure 1 and Table 1).
@@
Table 1: Spearman correlation between model size and psychological metrics. $\uparrow$ indicates positive correlation with size, $\downarrow$ indicates negative correlation. Significance: ***p<<0.001, **p<<0.01, *p<<0.05, n.s. = not significant
Although perplexity did not correlate significantly with scale, we found that perplexity correlated with question-level SD (Spearman’s $\rho$ = 0.465). This indicates that uncertainty partially, but incompletely, explains variability.
@@
The persona comparison analysis (Figure 1, right panel) shows that psychological trait expression is highly malleable through persona prompting, with effects that vary systematically by both trait and persona type.
Wilcoxon signed-rank tests reveal that positive personas (buddhist, teacher) exhibit significantly lower scores on negative traits (p<<0.001). The teacher persona also shows higher scores on positive traits (p<<0.001), while the buddhist persona shows significantly lower response variability and perplexity across model sizes (p<<0.05).
The clinical personas (antisocial, schizophrenia) show pronounced deviations from the baseline assistant persona, with lower scores for positive traits and elevated scores for negative traits (p<<0.001). For the schizophrenia persona, we also observe increased response variability and perplexity (p<<0.05).
A more detailed description of the results can be found in Tables A2-A5 in the Appendix .
Effects of Reasoning
Figure 2: Mean question-level variability (SD) and perplexity across Reasoning Effort levels (GPT-OSS) and Reasoning Mode On versus Reasoning Mode Off (Qwen-3, Qwen-3 MoE, DeepSeek, Claude). The analysis combines BFI and SD3 (71 items total) for the assistant persona with question re-ordering. Question-level variability tends to increase with reasoning effort and for reasoning versus non-reasoning models, while perplexity decreased for most models of the GPT-OSS and Qwen-3 families.
Figure 2 presents our results on chain-of-thought reasoning. Mean question-level variability increased with greater reasoning effort (GPT-OSS: aggregated Kruskal–Wallis test, p<<0.001; Dunn’s post hoc, all p<<0.05) and was higher in reasoning models relative to non-reasoning versions (Qwen-3, Qwen-3 MoE, DeepSeek, Claude: aggregated Mann–Whitney U tests, all p<<0.01). Perplexity, by contrast, generally decreased with reasoning (GPT-OSS: aggregated Kruskal–Wallis test, p<<0.001; Dunn’s post hoc, all p<<0.001; Qwen-3, Qwen-3 MoE: aggregated Mann–Whitney U test, all p<<0.001; DeepSeek: n.s.).
@@
Figure 3 compares the variability in responses between the original BFI and SD3 questionnaires and the LLM-adapted versions. This comparison reveals that the LLM-adapted instruments show similar question-level variability. Perplexity increased for the LLM-adapted questionnaires (Table 2). These findings suggest that the observed instability is not due only to the inappropriate application of human-centric concepts to LLMs.
@@
Figure 4 compares the effects of paraphrasing versus reordering statements (shuffling). We observed a trend towards a correlation between model size and the effect of paraphrasing on response variability (Spearman’s $\rho$ = 0.39, p = 0.0671). For larger models (>> 50B), paraphrasing significantly increased response variability (Table 3).
@@
Models << 50B (n=19) 0.244 n.s
Models $\geq$ 50B (n=4) <<0.01** $\uparrow$
@@
Figure 5 compares evaluations conducted with conversation history (i.e., presenting the questionnaire in a multi-turn format) to those conducted without conversation history. Larger models tend to show a decrease in question-level variability (SD) when conversation history is provided (Spearman’s $\rho$: -0.512; p-value: 0.0126), indicating increased consistency. However, Wilcoxon signed-rank tests show that while larger models (>> 50B) become more consistent, for smaller models (<< 50B) providing conversation history greatly increases their variability (Table 4).